OpenStudy (anonymous):

Use partial fractions: ∫(x−7)/(x^2−x−12)dx

OpenStudy (anonymous):

I got AX +3A +BX -4B = X-7 I solved and got A=-3/7 and B =10/7 which gave me \[\frac{ -3 }{ 7 }\ln(x-4)+\frac{ 10 }{ 7 } \ln(x+3)\] Anyone know where I went wrong?

OpenStudy (loser66):

why do you think it is wrong?

hartnn (hartnn):

you just missed the "+c" part after integration, everything else is correct.

OpenStudy (anonymous):

webworks says I got it incorrect, I attached a screenshot. Could you guys check to see if it looks right?

hartnn (hartnn):

please add the "+c" in your answer and check again

hartnn (hartnn):

i see its already there...

OpenStudy (anonymous):

If I add the +c, I get "Variable 'C' is not defined in this context"

OpenStudy (loser66):

you have the option "email instructor" :) I don't see anything wrong.

hartnn (hartnn):

ln(4-x) instead of ln(x-4)

OpenStudy (anonymous):

Thank you. I feel stupid

OpenStudy (loser66):

@ryanmk54 I am with you. :) don't know why ln (4-x)

OpenStudy (loser66):

If you say ln |x-4| , I am ok

OpenStudy (anonymous):

I see what you are saying now. Absolute value of x-4 = Absolute value -x+4